OpenStudy (anonymous):

find equation of tangent to curve y=2/x2-3 at given point (3,1/3|)

OpenStudy (anonymous):

First rewrited your function so that the power comes to the top. y=2x^-2 -3

OpenStudy (anonymous):

Then find the derivative so that you can determine the slope when x=3 dy/dx=2(-2)x^-3=-4/x^3

OpenStudy (anonymous):

Next substitute your x value to find the slope at that point -4/(3)^4=-4/81

OpenStudy (anonymous):

Then use point slope form to write the equation of the line y-1/3=(-4/81)(x-3)